Grade 11Mechanics

What will be the minimum velocity of vehicle on a circular banked road?

Priya
6 Years agoGrade 11

1 Answer

Arun
6 Years ago
 

Let us consider the situation given in the figure. Let r is the radius of turn which you have to take. N1 and N2 are normal reactions mg the weight and F1, F2 the frictional forces on the inside and outside wheels.

So, (N1 + N2) = mg                          ......... (i)

Thus, F1 + F2 = μ(N1 + N2) = μgm         ......... (ii)

This must be greater than or equal to the centripetal force required.

Therefore, μmg > mv2/r

or, v 

or, Vmax = √μrg